# Useful entanglement can be extracted from noisy graph states

### v1

@@ 5.1 Noise-correcting structures and robust families of states
As introduced above, noise processes lead to an ensemble of different states $\{(p_{i},\ket{\psi_{i}}\}$ being prepared, where one of the ensemble components is the ideal state |$\ket{G}$. This ensemble can be interpreted as a probabilistic preparation of an unknown state |$\ket{\psi_{i}}$ with probability $p_{i}$. The same measurement procedure as previously might now yield results inconsistent with |$\ket{G}$, and observations of those are an indication of state other than |$\ket{G}$ being prepared.
@@
As an example, consider the graph $G$ to be a $4$-cycle of vertices $\{1,2,3,4\}$, as shown in Figure 1a. The operator $O=X_{1}X_{3}$ belongs to a stabilizer of |$\ket{G}$, thus the only sign structures that can be observed in the sequential measurement of qubits $1$ and $3$ are $+1,+1$ and $-1,-1$. Consider now all edge subgraphs $\{G^{\prime}\}$ of $G$ and a physical process in which an ensemble of graph states $\{p_{G^{\prime}},\ket{G^{\prime}}\}$ is prepared, some of them are shown in Figure 1b. The operator $O$ (and $-O$) is not a part of the stabilizer of any proper subgraph with at least one edge; thus, sign structures of $+1,-1$ or $-1,+1$ are possible: observing those in the same measurement procedure implies that the pre-measurement state was not |$\ket{G}$.
@@
Figure 4: Considered families of graphs for the task of Bell pair extraction, parameterized by the length of the internal part $n$. The terminal qubits, designated for creation of a Bell pair, are denotes with filled nodes. From the top: path graph, twisted pair, crazy graph.
@@
The postselection on the right measurement outcomes consistent thus performs a probabilistic check whether |$\ket{G}$ was truly prepared. However, the existence of such parity-checking structures depends on the graph and the task for which the associated quantum state is used.
@@
A Bell pair can be prepared with the help of a path graph state and local $X$ measurements, but such a system does not support embedded stabilizer operators. Other graphs and measurement patterns do better in this regard, and here we present them. The families of graph states analyzed are parameterized by the length between the terminal qubits across which a Bell pair is to be prepared (see Figure 4). Apart from a simple path graph, included as a benchmark for comparison, we include two other families described below – both of them have their strengths.
@@
The twisted pair graph [37] is built from layers, where each qubit in the layer $k$ is connected with every qubit from the layers $k-1$ and $k+1)$. The number of qubits in each layer alternates between $1$ and $2$, which enables the existence of a restricted set of stabilizers: for each 2-qubit layer consisting of the qubits $\{i,j\}$, the operators $X_{i}X_{j}$ stabilize |$\ket{G}$. Additionally, it is very robust to certain types of noise, because of its relatively simple geometry.
@@
The crazy graph (studied in [38, 14] for its noise robustness) also consists of layers of $2$ qubits each, with a similar full connectivity structure between the adjacent layers. This ensures a robust and simple to analyze structure of the embedded stabilizer operators (see Appendix A and Figure 5).
@@
Both presented structures generalize to the extraction of other types of states, in which context they serve as a direct replacement of path graphs. We mention the usage in a task of 3-qubit GHZ state extraction.
5.2 Bell pair quality scaling for different type of noise
@@
Each of the graph states associated with the aforementioned three families (path graph, twist graph, and crazy graph) can be used for extraction of a Bell pair across the terminal qubits. Only local $X$ measurements are used for all three, and the fact that in the ideal case a maximally entangled state is produced is captured by the stabilizer operators consistent with this measurement pattern.
@@
As a simple example, consider the 5 qubit line graph as discussed in Section 3. Measurements in the $X$ basis on all but the terminal qubits lead to a two-qubit state which can be shown to have Bell-like correlations. Similar structures can be found for all the three graph families mentioned: sets of stabilizer operators ensuring correlations of the terminal qubits, having only $X$ operators in the internal section (and thus compatible with the measurement scheme).
The effect of noise destroys the perfect correlations: even if for every ensemble element the end state is maximally entangled, the mixture of such states might have less (or none) entanglement. Postselection on the measurement outcomes consistent with the ideal stabilizer structure helps twofold: one of its effects is probabilistic error detection. The other, appearing in the case of the crazy graph ensemble, is more subtle: after postselection, the end state may be the same for a large portion of the ensemble elements (see Appendix A).
@@
The results, presented in Figure 6, show that this approach (noise-detecting structures combined with postselection) works for different types of noise. The fidelity susceptibility (quantifying the initial fidelity falloff in the low noise limit; see Equation 24) is reduced for the twisted pair and crazy structures. Furthermore, the susceptibility for the crazy graph structure does not depend on the length, as a result of the correctional stabilizers mentioned in the previous paragraph. This behavior is observed both for uncorrelated phase noise, resulting in probabilistic edge losses and a local $Z$ flip noise.
An additional entanglement-preserving structure can be found in the case of perfectly correlated phase noise. The authors of Ref. [13] recently observed that in the case of the path graph, postselection on different measurement outcomes yields vastly different results in the terms of entanglement quality; we have found that this result does generalize to the other linear graphs analyzed by us. Both for the crazy and double twist graph families the lowest order nontrivial noise effects can be canceled completely if all the measurement outcomes are postselected on measuring the minus outcome of $X$. This is consistent with our previous observations: such an observed measurement pattern does not violate parity constraints arising from the stabilizer structure. This result can be understood by studying the modifications of the stabilizer operators of |$\ket{G}$ under correlated controlled-phase noise.
For the square graph presented in Figure 1a, if instead of the perfect controlled-$Z$ unitary, the $\operatorname{C\Phi}(\pi+\varepsilon)=\operatorname{diag}(1,1,1,-1\exp(i\varepsilon))$ is used (so that $\operatorname{C\Phi}=\operatorname{CZ}$ for $\varepsilon=0$), the four-qubit entangled state is a weighted graph state |$\ket{G}_{\varepsilon}$. If qubits 1 and 3 are measured in the $X$ basis, we expect the measurement outcomes to be equal (we exclude the cases when it does not hold). With the weighted graph state |$\ket{G}_{\varepsilon}$ as the initial state, let us denote the unnormalized post-measurement state by
|$\ket{X_{1,3}\pm}=\frac{\mathbbm{1}\pm X_{1}}{2}\frac{\mathbbm{1}\pm X_{3}}{2}\ket{G}_{\varepsilon}.$ (25)
The relevant expectation values defining the end correlations (with $\langle O\rangle_{\pm}$ denoting the expectation value of $O$ for the unnormalized state |$\ket{X_{1,3}\pm}$) across qubits $2$ and $4$ are the following, in the lowest nontrivial noise order:
$\langle Z_{2}Z_{4}\rangle_{+}$ $=1-\varepsilon^{2},$ $\langle Z_{2}Z_{4}\rangle_{-}$ $=-1+\frac{\varepsilon^{2}}{2},$ (26)
$\langle X_{2}X_{4}\rangle_{+}$ $=1-3\varepsilon^{2},$ $\langle X_{2}X_{4}\rangle_{-}$ $=-1+\frac{\varepsilon^{2}}{2},$
$\langle Y_{2}Y_{4}\rangle_{+}$ $=-1+3\varepsilon^{2},$ $\langle Y_{2}Y_{4}\rangle_{-}$ $=1+\frac{\varepsilon^{2}}{2},$
$\||X_{1,3}+\rangle\|^{2}$ $=1-\varepsilon^{2},$ $\||X_{1,3}-\rangle\|^{2}$ $=1-\frac{\varepsilon^{2}}{2}.$
Note that for the |$\ket{X_{1,3}-}$ outcome the squared length $\||X_{1,3}-\rangle\|^{2}$ is proportional to the unnormalized expectation values: thus, for the physical correlations the dominant noise effects are canceled and only higher order terms ($O(\varepsilon^{4})$) remain. On the other hand, if the outcome is |$\ket{X_{1,3}+}$, the noise is amplified compared to postselection on only stabilizer-consistent outcomes. Thus, further postselection can amplify the entanglement quality of the remaining qubits at the cost of discarding certain outcomes and reduced production rates.
The double twist graph can be viewed as multiple such squares stacked together by the terminal vertices: thus, a proper postselection can mitigate the first nontrivial noise effects completely (see Figure 7). A similar structure appears for the crazy graph: it is much more complex to analyze it, but computer algebra systems provide evidence of a periodic structure of susceptibilities as the length of the graph increases.
@@
Figure 6: Susceptibility of the Bell pair entanglement quality to edge loss (top) and local phase flip (bottom) noise. Pictured is the linear part of concurrence decrease for the three analyzed families (with postselection on the measurement outcomes consistent with the ideal state) in the low noise limit, where the dominant contributions are the single edge losses (top) and single qubit flips (bottom). The constant susceptibility of the crazy graph is a result of the noise at the terminal edges only: each internal edge loss is automatically corrected for and every single flipped qubit is detected.
@@
Figure 7: Susceptibility of the Bell pair entanglement quality for correlated phase noise. Here, only the cases postselected on observing exclusively minus signs during the $X$ measurement are taken into account. In this regime, the twist graph is able to remove the effect of noise completely (in the lowest nontrivial orders), and the crazy graph shows a periodic pattern of susceptibilities, still significantly reducing the noise effect. The situation inverts for observing exclusively $+$ signs: then, the noise gets amplified.
6 Extraction of GHZ states
The methodology applied for Bell pair extraction generalizes to the extraction of GHZ states [23]: generally, for $N$ qubits, they correspond to a star graph with a central qubit connected to the remaining $N-1$ ones (or, equivalently, a complete graph of $N$ qubits). Here, we concentrate on the case $N=3$ and adopt star-shaped graph structures built from the graph patterns used in Bell pair creation. Refer to Figure 8a for an illustration of the structure: it is built on the template of several crazy graphs, stitched at the central pair of vertices. The fidelity response, presented in Figure 8b, is more resilient to noise (as compared to the simpler path graph) also in this new task of extraction of the GHZ state.
The data presented in this figure require calculations involving a large number of ensemble states and measurement outcomes. To mitigate the numerical difficulty, a Monte Carlo method was utilized. The figure of merit – fidelity – is linear with respect to the ensemble- and measurement outcome averaging, provided the target state is pure. Thus, this approach allows for an approximation of fidelity by sampling ensemble elements and measurement outcomes, avoiding direct calculations with density operators. The fidelity of the output state to the ideal pure state is estimated using this method until the plots converge smoothly and the variance of the fidelity estimator is negligible.
Analysis of the data in Figure 8 shows that the fidelity response for both the GHZ and Bell state extractions exhibits qualitatively similar behavior: most importantly, the susceptibility is greatly reduced in the low noise limit. Thus, the ’crazy graph’ structures can generalize beyond Bell pair extraction to more complex multi-qubit systems, maintaining their noise-resilient properties. This performance gain is due to the noise-detecting and -correcting capabilities intrinsic to these graph families. Further investigations of different connectivity structures suggest that this behavior is universal, although a formal proof is beyond the scope of this study. If true, the ’crazy graph’ can serve as a link similar to a simple path link, similar in function but implementing some level of error correction.
@@ 6 Extraction of GHZ states
Figure 8: a) Structures of the analyzed graphs analyzed: path and crazy templates yielding Bell pairs and 3-qubit GHZ states upon measurements of the central qubits. b) Fidelity of the end state with respect to the ideal Bell and GHZ states as a function of noise. The fidelity is averaged over possible measurement outcomes consistent with the measurement-embedded stabilizer structures. The initial fall-off rate of fidelity at $p=0$ is the fidelity susceptibility, defined in Equation 24.
@@ 7 Summary and Outlook
This study has advanced the understanding of noise processes in graph states, specifically addressing the effects of preparation by the Ising-like interaction and photonic qubit fusion. We have found graph structures applicable in error detection and correction, especially in tasks such as information transfer and GHZ state preparation. This approach uses additional qubits for probabilistic verification and demonstrates a new method to improve resilience against specific noise disturbances.
Future research may include more general tasks found in the measurement-based quantum computation approach. So far, our results are applicable to measurement patterns composed of Pauli operators, but it is known to restrict the set of possible operations to a clasically simulable one: a similar method, based on non-Pauli measurements, but still allowing for noise effects reduction, could open new pathways for quantum computation, albeit with conceptual and practical complexities. The methods presented here can be generalized for arbitrary output graph geometry, and the applicability of this procedure for quantum metrology [39] could be studied.
@@ Acknowledgments
We thank Frederik Hahn, Mariami Gachechiladze, and Jan L. Bönsel for discussions. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project numbers 447948357 and 440958198), the Sino-German Center for Research Promotion (Project M-0294), the ERC (Consolidator Grant 683107/TempoQ), the German Ministry of Education and Research (Project QuKuK, BMBF Grant No. 16KIS1618K), the Stiftung der Deutschen Wirtschaft, and the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement (No 945422).
Appendix A Post-measurement correlations and stabilizers compatible with the measurement scheme

### v2

@@ 5.1 Noise-correcting structures and robust families of states
As introduced in Section 4, noise processes lead to an ensemble of different states $\{(p_{i},\ket{\psi_{i}})\}$ being prepared, where one of the ensemble components is the ideal state |$\ket{G}$. This ensemble can be interpreted as a probabilistic preparation of an unknown state |$\ket{\psi_{i}}$ with probability $p_{i}$. The measurement procedure might now yield results inconsistent with |$\ket{G}$, and observations of those are an indication of state other than |$\ket{G}$ being prepared – see Corollary 2.
@@
As an example, consider the square graph – a $4$-cycle of vertices $\{1,2,3,4\}$, as shown in Figure 2a. The operator $O=X_{1}X_{3}$ is an element of the stabilizer of $\ket{\square}$, thus the only sign structures that can be observed in the sequential measurement of qubits $1$ and $3$ are $+1,+1$ and $-1,-1$. Consider now all edge subgraphs $\{G^{\prime}\}$ of the square graph and a physical process in which an ensemble of graph states $\{p_{G^{\prime}},\ket{G^{\prime}}\}$ is prepared, some of them are shown in Figure 2b. The operator $O$ (and $-O$) is not a part of the stabilizer of any proper subgraph with at least one edge; thus, sign structures of $+1,-1$ or $-1,+1$ are possible. Observing those in the same measurement procedure implies that the pre-measurement state was not $\ket{\square}$.
@@
Figure 4: Considered families of graphs for the task of Bell pair extraction, parameterized by the length of the internal part $n$. The terminal qubits, designated for creation of a Bell pair, are denoted with filled nodes. From the top: path graph, twisted pair, crazy graph.
@@
Therefore, the postselection on the right measurement outcomes performs a probabilistic check whether |$\ket{G}$ was truly prepared. However, the existence of such parity-checking structures depend on the graph and the task for which the associated quantum state is used.
@@
A Bell pair can be prepared with the help of a path graph state and local $X$ measurements (Figure 3a), but such a system does not support embedded stabilizer operators. Other graphs and measurement patterns do better in this regard. Here we present some alternatives.
The families of graph states analyzed are parameterized by the length between the terminal qubits across which a Bell pair is to be prepared (see Figure 4). Apart from a simple path graph, included as a benchmark for comparison, we include two other families described below – both of them have their strengths.
@@
The twisted pair graph [50], shown in Figure 4b, is built from layers, where each qubit in the layer $k$ is connected with every qubit from the layers $k-1$ and $k+1$. The number of qubits in each layer alternates between $1$ and $2$, which enables the existence of a restricted set of stabilizers: for each 2-qubit layer consisting of the qubits $\{i,j\}$, the operators $X_{i}X_{j}$ stabilize |$\ket{G}$. Additionally, it is very robust to certain types of noise, because of its relatively simple geometry.
@@
The crazy graph similarly consists of layers of $2$ qubits each, with full connectivity between the adjacent layers. It is known for its noise robustness [51, 24], which stems from the structure of its stabilizer operators. See Figure 5 and Appendix B for more details.
@@
Each of the graph states associated with the aforementioned three families (path graph, twist graph, and crazy graph) can be used for extraction of a Bell pair across the terminal qubits (shaded vertices in Figure 4). Only local $X$ measurements are used for all three, and in the ideal case a maximally entangled state is produced. This is because there exist stabilizer operators consistent with this measurement pattern, as in the exemplary 4 qubit path graph discussed in Section 3.
@@
Similar structures can be found for all the three graph families mentioned. We find sets of stabilizer operators which have $X$ operators in the internal section and Pauli operators which ensure correlations on the terminal qubits – see Section B.2. These operators are compatible with the measurement scheme.
5.2 Bell pair quality scaling for different types of noise
@@
We analyze the approach of noise-detecting structures for graph families introduced above. The procedure is employed for different layer sizes $n$, see Figures 6 and 7. We find that the twisted pair and crazy structures are less affected by the noise, as evidenced by the fidelity susceptibility $\alpha$. The value of $\alpha$ is smaller, and in the case of the crazy graph does not depend on the graph length. Furthermore, the susceptibility for the crazy graph structure does not depend on the length, as a result of the correctional stabilizers mentioned in the previous paragraph. This behavior is observed both for uncorrelated phase noise, resulting in probabilistic edge losses and a local $Z$ flip noise.
The results can be understood as destruction of the perfect correlations present in graph states, mathematically described in Section 3. Even if for every ensemble element the end state is maximally entangled, the mixture of such states might have less or no entanglement. Postselection on the measurement outcomes consistent with the ideal stabilizer structure helps twofold: one of its effects is probabilistic error detection. The other, appearing in the case of the crazy graph ensemble, is more subtle: after postselection, the end state may be the same for a large portion of the ensemble elements (see Appendix B).
@@
Figure 6: The effects of the edge loss (top) and local phase flip (bottom) noise on the task of Bell pair extraction for the three analyzed graph state families. The susceptibility $\alpha$ (Equation 22) is the initial falloff of fidelity in the low noise limit of the resulting two-qubit state with respect to the expected Bell pair. The results here only take into account the measurement outcomes consistent with the ideal state. In this case (low noise limit) the dominant contributions are the single edge losses (top) and single qubit flips (bottom). The constant susceptibility of the crazy graph is a result of the noise at the terminal edges only: each internal edge loss is automatically corrected for and every single flipped qubit is detected.
An additional entanglement-preserving structure can be found in the case of perfectly correlated phase noise. This is the source of zero susceptibility of twist graph and periodic behavior of crazy graph in Figure 7. The authors of Reference [21] recently observed that in the case of the path graph, postselection on different measurement outcomes yields vastly different results in the terms of entanglement quality. We have found that this result does generalize to the other linear graphs analyzed by us. Both for the crazy and twist graph families the lowest order nontrivial noise effects can be canceled completely if all the measurement outcomes are postselected on measuring the $(-1)$ outcome of $X$.
@@
Figure 7: The effects of correlated phase noise on the task of Bell pair extraction for the three analyzed graph state families. The susceptibility $\alpha$ (Equation 22) is the initial falloff of fidelity in the low noise limit of the resulting two-qubit state with respect to the expected Bell pair. Here, only the cases postselected on observing exclusively minus signs during the $X$ measurement are taken into account. In this case (low noise limit with restrictive postselection) the twist graph is able to remove the effect of noise completely, and the crazy graph shows a periodic pattern of susceptibilities, still significantly reducing the noise effect. The situation inverts for observing exclusively $+$ signs: then, the noise gets amplified.
This is consistent with our previous observations: such an observed measurement pattern does not violate parity constraints arising from the stabilizer structure. This result can be understood by studying the modifications of the stabilizer operators of |$\ket{G}$ under correlated controlled-phase noise – see Appendix D.
6 Extraction of GHZ states and large noise regime
@@
Figure 8: a) Structures of the analyzed graphs analyzed: path and crazy templates yielding Bell pairs and 3-qubit GHZ states upon measurements of the central qubits. Similarly to the previous case, all qubits except the terminal ones (highlighted and enlarged) are measured in the $X$ basis. b) Fidelity of the end state with respect to the ideal Bell and GHZ states as a function of noise. The fidelity is averaged over possible measurement outcomes consistent with the measurement-embedded stabilizer structures. The initial fall-off rate of fidelity at $p=0$ is the fidelity susceptibility, defined in Equation 22.
In Section 5 we concentrated on the Bell pair extraction for various geometries in the low noise regimes. However, the presented structures (twisted and crazy graphs) are of more general use. In this section, we analyze large noise effects on two exemplary scenarios. In addition to Bell pair extraction, we consider a similar procedure for GHZ states [35] arising from a starlike graph. The considered graph structures are presented in Figure 8a. The main result here is the fidelity of output states (Figure 8b) postselected on observing the expected sign structures during $X$ measurement of the internal qubits.
The Monte Carlo simulations required to analyze the large noise regime are lengthy, as multiple ensemble elements are relevant for the result. Therefore, for simplicity we consider fixed graph geometries (path and crazy-like with fixed length) with only the uncorrelated phase noise present.
The fidelity $\mathcal{F}$ is determined as the average of ensemble elements and postselected stabilizer-consistent outcomes – see Equation 21. This avoids direct calculations with density operators. The fidelity of the output state to the ideal pure state is estimated using this method until the plots converge smoothly and the variance of the fidelity estimator is negligible.
Analysis of the Figure 8 shows that the fidelity response for both the GHZ and Bell state extractions exhibits qualitatively similar behavior. The output state is more resilient to noise (as compared to the simpler path graph) also in this new task of extraction of the GHZ state. Most importantly, the noise effect is greatly reduced in the intermediate noise limit. This performance gain is due to the postselection involving noise-detecting and -correcting capabilities intrinsic to this graph family. Thus, the ‘crazy graph’ structures can serve as a general link. They can generalize beyond Bell pair extraction to more complex multi-qubit systems, maintaining their noise resistant properties.
Further investigations of different connectivity structures suggest that this behavior is universal, although a formal proof is beyond the scope of this study. If true, the ‘crazy graph’ can serve as a link similar to a simple path link. That is, it has similar in functions but some level of error correction is implemented.
@@ 7 Summary and Outlook
This study has advanced the understanding of noise processes in graph states, specifically addressing the effects of preparation by the Ising-like interaction and photonic qubit fusion. We have found graph structures applicable in error detection and correction, especially in tasks such as information transfer and GHZ state preparation. This approach uses additional qubits for probabilistic verification and demonstrates a new method to improve resilience against specific noise disturbances. The mathematical formalism used in the calculations has been previously used in the nonlocality detection in graph states [52, 53] and could be extended to other families beyond Bell and GHZ states.
Future research may include more general tasks found in the measurement-based quantum computation approach. So far, our results are applicable to measurement patterns composed of Pauli operators, but it is known to restrict the set of possible operations to a classically simulable one. A similar method, based on non-Pauli measurements, but still allowing for noise effects reduction, could open new pathways for quantum computation, albeit with conceptual and practical complexities. The methods presented here can be generalized for arbitrary output graph geometry, and the applicability of this procedure for quantum metrology [54] could be studied.
@@ Acknowledgments
We thank Frederik Hahn, Mariami Gachechiladze, and Jan L. Bönsel for discussions. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project numbers 447948357 and 440958198), the Sino-German Center for Research Promotion (Project M-0294), the ERC (Consolidator Grant 683107/TempoQ), the German Ministry of Education and Research (Project QuKuK, BMBF Grant No. 16KIS1618K), the Stiftung der Deutschen Wirtschaft, and the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement (No 945422). Funding from the projects DeQHOST APVV-22-0570, QUAS VEGA 2/0164/25, and Stefan Schwarz programme is gratefully acknowledged.
Appendix A Bell pair creation is equivalent to teleportation
As mentioned in the introduction, the creation of a Bell pair is equivalent to the task of information transfer (teleportation) across the graph. This is so because if the Bell pair creation in the graph $G$ is viewed in the context of a larger graph $G^{\prime}$, the entire process, followed by measurement of one of the terminal qubits, performs qubit fusion: the unmeasured terminal qubit acts as the two combined. Since all local measurements with nonoverlapping supports do commute, this can be done before any other measurement or afterwards: in the latter case, the information created in one of the qubits is merged with the second one. This is captured by the following Lemma, represented pictorially by the Figure 9.